Circular No. 8314

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2004 CB (LINEAR) An apparently asteroidal object discovered by the LINEAR project

(discovery observation below from MPEC 2004-C16 and MPS 98921) has been found to show a narrow 1′.1 tail in p.a. 274^{o} (slightly expanding toward the end) on CCD images obtained by R. H. McNaught with the 1.0-m f/8 reflector at Siding Spring on Mar. 30.8 UT. Following a request by the Central Bureau, M. Kočer reports that CCD frames taken at Kleť on Mar. 31.145 also show a narrow tail ∼ 90′′ long in p.a. ≈ 280^{o}.

2004 UT  α_{2000}  δ_{2000}  Mag.
Feb. 3.40369  14^{h}24^{m}23.^{s}21  +43^{o}34′21′′.8  18.1

McNaught’s astrometry, the following orbital elements, and an ephemeris appear on MPEC 2004-F96.

Epoch = 2004 Mar. 16.0 TT

T = 2004 Apr. 2.1633 TT  ω = 149.^{o}6558
e = 0.689400  Ω = 66.4875 }2000.0
q = 0.912180 AU  i = 19.1472

a = 2.936829 AU n^{o} = 0.1958331 P = 5.033 years

SUPERNOVAE 2004ba AND 2004bb D. Norman, Cerro Tololo Interamerican Observatory (CTIO); G.

Squires, Space Infrared Telescope Facility, California Institute of Technology; J. Kubo, Brown University; and A. Becker, University of Washington, on behalf of the ‘Deep Lens Survey’ team, report the detection of three possible supernovae in R and (subsequent) V -band images taken with the CTIO 4-m reflector (+ MOSAIC imager). The two objects with apparent hosts are tabulated below.

SN  2004 UT  α_{2000}  δ_{2000}  R  Offset
2004ba Mar. 20.299 13^{h}56^{m}56.^{s}54 −11^{o}25′39′′.9 22.9 0′′.6 W, 1′′.6 N
2004bb Mar. 20.240 13 57 10.47 −11 03 19.5 21.7 4′′.0 E, 0′′.7 N

Additional magnitudes: 2004ba, 2003 Apr. 2 UT, R [25.3; 2004 Mar. 23.255, V = 22.4. 2004bb, 2003 Apr. 2, R [25.3; 2004 Jan. 18, V [25.4; Mar. 24.181, V = 22.1. The third transient, with no visible galaxy nearby, is located at α = 13^{h}55^{m}31^{s}.91, δ = −11^{o}50′57′′.2; available magnitudes: 2003 Apr. 2, R [25.3; 2004 Mar. 20.299, R = 23.0; 23.255, V = 23.0.

2004 March 31  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green